solve 2x^2-3x-14=0
Answer by oberobic(2304)

The first approach is to try to factor it.
You know you will have: (2x )(x ) for the x terms. And since the constant is negative, one of the factors will have a - and the other will have a +.
You know that 14 = 2 * 7, which are your only choices.
So the only choices are
(2x + 2)(x - 7)
(2x - 2)(x + 7)
(2x + 7)(x - 2)
(2x - 7)(x + 2)
You're looking to the -3x, which is:'
(2x - 7)(x + 2)
ALWAYS check!
(2x - 7)(x + 2) = 2x*x + 2x*2 - 7*x - 7*2 = 2x^2 + 4x - 7x - 14 = 2x^2 - 3x - 14.